# Descent Methods for Vector Optimization Problems: A Majorization-Minimization Perspective

### v1

@@ Abstract
In this paper, we develop a unified framework and convergence analysis of descent methods for vector optimization problems (VOPs) from a majorization-minimization perspective. By choosing different surrogate functions, the generic method reduces to some existing descent methods with and without line search, respectively. The unified convergence analysis shows that the slow convergence of the steepest descent method is mainly due to the large gap between surrogate and objective functions. As a result, the performance of descent methods can be improved by narrowing the surrogate gap. Interestingly, we observe that selecting a tighter surrogate function is equivalent to using an appropriate base of the dual cone in the direction-finding subproblem. Furthermore, we use Barzilai-Borwein method to narrow the surrogate gap and devise a Barzilai-Borwein descent method for VOPs with polyhedral cone. By reformulating the subproblem, we provide a new insight into the Barzilai-Borwein descent method and bridge it to the steepest descent method. Finally, several numerical experiments confirm the efficiency of the proposed method.
@@ 1 Introduction
In vector optimization, a vector-valued function $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is to be optimized under the partial order induced by a closed, convex, and pointed cone $K\subset\mathbb{R}^{m}$ with a non-empty interior, defined as follows:
@@
Let $K^{*}=\{c^{*}\in\mathbb{R}^{m}:\langle c^{*},y\rangle\geq 0,\forall y\in K\}$ be the positive polar cone of $K$, and $C$ be a compact convex set such that $0\notin C$ and $\mathrm{cone}(C)=K^{*}$, namely, $C$ is a base of $K^{*}$. This paper focuses on the following vector optimization problem:
${\rm min}_{K}~{}F(x),$ (VOP)
where $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is differentiable. In vector optimization, it is often impossible to improve all objectives simultaneously with respect to the partial order. Therefore, the concept of optimality is defined as efficiency [24], meaning that there is no better solution for an efficient solution. Specifically, the problem (VOP) corresponds to a multiobjective optimization problem when $K=\mathbb{R}^{m}_{+}$, where +$\mathbb{R}^{m}_{+}$ denotes the non-negative orthant of $\mathbb{R}^{m}$. Various applications of multiobjective optimization problems (MOPs) can be found in engineering [30], economics [40, 19], management science [14], environmental analysis [27], machine learning [36, 42], etc. Although many real-world problems reformulated as vector-valued problems adhere to the partial order induced by +$\mathbb{R}^{m}_{+}$, some applications, such as portfolio selection in securities markets [1, 2], require partial orders induced by closed convex cones other than the non-negative orthant. Consequently, vector optimization problems (VOPs) have garnered significant attention in recent years.
Over the past two decades, descent methods have received increasing attention within the multiobjective optimization community, primarily due to the seminal work on the steepest descent method proposed by Fliege & Svaiter [16]. Inspired by Fliege and Svaiter’s contributions, researchers have extended other numerical algorithms to solve multiobjective optimization problems (MOPs) [see, e.g., 15, 35, 33, 17, 5, 31, 32, 38]. To the best of our knowledge, the study of descent methods for unconstrained vector optimization problems can be traced back to the work of Graña Drummond & Svaiter [13], who extended the steepest descent method for MOPs (SDMO) proposed by Fliege & Svaiter [16] to VOPs. In this context, the direction-finding subproblem at $x^{k}$ is formulated as follows:
@@
In recent years, complexity analysis of descent methods for MOPs has been extensively studied. Fliege et al. [18] established the convergence rates of SDMO under different convexity assumptions. Similarly, Tanabe et al. [39] developed convergence results for the multiobjective proximal gradient method. Chen et al. [8] noted that both theoretical and empirical results indicate that multiobjective first-order methods exhibit slow convergence due to objective imbalances. To address this challenge, Chen et al. [7] proposed a Barzilai-Borwein descent method for MOPs (BBDMO) that dynamically tunes gradient magnitudes using Barzilai-Borwein’s rule [3] in direction-finding subproblem. Moreover, an improved linear convergence rate is confirmed by Chen et al. [8], demonstrating that Barzilai-Borwein descent methods can effectively mitigate objective imbalances from a theoretical perspective. Despite the extensive study of the complexity analysis of descent methods for MOPs, corresponding results for VOPs have received little attention. As described by Chen et al. [6], the linear convergence rates of first-order descent methods for VOPs are influenced by $C$, the base of $K^{*}$ in direction-finding subproblem. This naturally leads to the following question:
@@
(Q2) Is $\Delta_{m}$ a good choice of base for +$\mathbb{R}^{m}_{+}$ in SDMO? If not, how to choose a better base?
@@
Motivated by majorization-minimization optimization [29], which involves the successive minimization of a sequence of upper bounds of the objective function, we provide a unified convergence analysis of descent methods for VOPs from a majorization-minimization perspective. We emphasize that the gap between the surrogate and objective functions significantly affects the performance of descent methods, which plays a central role in majorization-minimization optimization. Specifically, we highlight that the steepest descent method for VOPs exhibits slow convergence due to the large gap between the surrogate and objective functions. To address this issue, we develop an improved descent method with a tighter surrogate function, resulting in improved linear convergence. Interestingly, we show that selecting a tighter surrogate function is equivalent to using an appropriate base in the direction-find subproblem. This provides a positive answer to (Q1). Additionally, we devise a Barzilai-Borwein descent method for VOPs (BBDVO) with polyhedral cones. By reformulating the subproblem, we observe that the BBDMO is essentially the SDMO with an appropriate base in direction-finding subproblem. This observation provides a positive answer to (Q2) and offers a new insight into BBDMO from a majorization-minimization perspective. Both theoretical and empirical results indicate that the performance of the proposed method is not sensitive to the choice of the transform matrix.
@@
The rest of this paper is organized as follows. Some notations and definitions are given in Section 2 for our later use. In Section 3, we propose a generic majorization-minimization descent method for VOPs and analyze its convergence rates under different convexity assumptions. Section 4 is devoted to showing the connections between different descent methods from a majorization-minimization perspective. In Section 5, the effect of the gradient adjustment is clarified from the vantage point of alleviating objective imbalances, and a Barzilai-Borwein descent method is devised for VOPs with polyhedral cone. Section 7 presents numerical results demonstrating the efficiency of the proposed Barzilai-Borwein method. Finally, conclusions are drawn at the end of the paper.
@@ 2.1 Vector optimization
By the differentiability of $F(\cdot)$, $K$-convexity of $F(\cdot)$ is equivalent to
@@ 2.2 Strong convexity and smoothness
Comparing with the definition in [12], Definition 2.6 includes the case $\bm{\mu}\in{\rm{bd}}(K)$, then it reduces $K$-convexity when $\bm{\mu}=0$.
Remark 2.3.
Assume that the objective function $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $x^{*}\in\mathbb{R}^{n}$ is a $K$-stationary point of (VOP) if and only if $x^{*}$ is an efficient solution of (VOP).
By the differentiability of $F(\cdot)$, strong $K$-convexity of $F(\cdot)$ is equivalent to
@@
In other words, the condition number of $F(\cdot)$ does not depend on $C$, but on $K$.
@@ 2       Choose a strongly $K$-convex surrogate function $G_{k}(\cdot)$ of $F(\cdot)-F(x^{k})$ near $x^{k}$
3       Update $x^{k+1}:=\mathop{\arg\min}_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x)}\right\rangle$
@@
    $F(x^{k+1})-F(x^{k})\preceq_{K}G_{k}(x^{k+1})$, where $x^{k+1}$ is the minimizer of $\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x)}\right\rangle$, furthermore, when $F(\cdot)-F(x^{k})\preceq_{K}G_{k}(\cdot)$ for all $x\in\mathbb{R}^{n}$, we call $G_{k}(\cdot)$ a majorizing surrogate;
@@
Let $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $x^{k+1}$ be the minimizer of $\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x)}\right\rangle$. Then, for all $x\in\mathbb{R}^{n}$, we have
@@
    $\left\langle{c^{*}_{k},F(x^{k+1})}\right\rangle+\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle\leq\left\langle{c^{*}_{k},F(x)}\right\rangle+\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle$, where $c^{*}_{k}$ is a maximizer of $\max_{c^{*}\in C}\min_{x\in\mathbb{R}^{n}}\left\langle{c^{*},G_{k}(x)}\right\rangle$.
@@
Assertion (i) directly follows by the $K$-smoothness of $H_{k}(\cdot)$ and the facts that $H_{k}(x^{k})=0$ and $JH_{k}(x^{k})=0$. Next, we prove the assertion (ii). By Sion’s minimax theorem [37], by denoting $c^{*}_{k}$ a maximizer of $\max_{c^{*}\in C}\min_{x\in\mathbb{R}^{n}}\left\langle{c^{*},G_{k}(x)}\right\rangle$, and $x^{k+1}$ the minimizer of $\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x)}\right\rangle$, we have $JG_{k}(x^{k+1})^{T}c^{*}_{k}=0$. This, together with the strong $K$-convexity of $G_{k}(\cdot)$, implies that

### v2

@@ Abstract
In this paper, we develop a unified framework and convergence analysis of first-order methods for vector optimization problems (VOPs) from a majorization-minimization perspective. By selecting different surrogate functions, the general method can be reduced to various existing first-order methods, both with and without line search. Our unified convergence analysis reveals that the slow convergence of the steepest descent method is primarily attributed to the significant gap between the surrogate and objective functions. Consequently, narrowing this surrogate gap can enhance the performance of first-order methods. Interestingly, we elucidate that selecting a tighter surrogate function is equivalent to using an appropriate base of the dual cone in the direction-finding subproblem. Building on this insight, we employ the Barzilai-Borwein method to narrow the surrogate gap and propose a Barzilai-Borwein descent method for VOPs (BBDVO) with polyhedral cones. By reformulating the subproblem, we provide a novel perspective on the Barzilai-Borwein descent method, bridging the gap between this method and the steepest descent method. Furthermore, by selecting an appropriate base and reformulating the subproblem in the Newton method, we also provide a novel perspective on the preconditioned Barzilai-Borwein method. Finally, several numerical experiments are presented to validate the efficiency of the BBDVO.
@@ 1 Introduction
This paper focuses on the following vector optimization problem:
${\rm min}_{K}~{}F(x),$ (VOP)
where $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is to be optimized under the partial order induced by a closed, convex, and pointed cone $K\subset\mathbb{R}^{m}$ with a non-empty interior, defined as follows:
@@
Let $K^{*}=\{c^{*}\in\mathbb{R}^{m}:\langle c^{*},y\rangle\geq 0,\forall y\in K\}$ be the positive polar cone of $K$, and $C$ be a compact convex set such that $0\notin C$ and $\mathrm{cone}(C)=K^{*}$, namely, $C$ is a base of $K^{*}$. In vector optimization, it is often impossible to improve all objectives simultaneously with respect to the partial order. Therefore, the concept of optimality is defined as efficiency [25], meaning that there is no better solution for an efficient solution. Specifically, the problem (VOP) corresponds to a multiobjective optimization problem when $K=\mathbb{R}^{m}_{+}$, where +$\mathbb{R}^{m}_{+}$ denotes the non-negative orthant of $\mathbb{R}^{m}$. Various applications of multiobjective optimization problems (MOPs) can be found in engineering [35], economics [45, 20], management science [15], environmental analysis [32], machine learning [41, 47], etc. Although many real-world problems reformulated as vector-valued problems adhere to the partial order induced by +$\mathbb{R}^{m}_{+}$, some applications, such as portfolio selection in securities markets [1, 2], require partial orders induced by closed convex cones other than the non-negative orthant. Consequently, vector optimization problems (VOPs) have garnered significant attention in recent years.
Over the past two decades, descent methods have received increasing attention within the multiobjective optimization community, primarily due to the seminal work on the steepest descent method proposed by Fliege & Svaiter [17]. Inspired by Fliege and Svaiter’s contributions, researchers have extended other numerical algorithms to solve multiobjective optimization problems (MOPs) [see, e.g., 16, 40, 38, 18, 5, 36, 37, 43]. To the best of our knowledge, the study of descent methods for unconstrained vector optimization problems can be traced back to the work of Graña Drummond & Svaiter [14], who extended the steepest descent method for MOPs (SDMO) proposed by Fliege & Svaiter [17] to VOPs. In this context, the direction-finding subproblem at $x^{k}$ is formulated as follows:
@@
In recent years, complexity analysis of descent methods for MOPs has been extensively studied. Fliege et al. [19] established the convergence rates of SDMO under different convexity assumptions. Tanabe et al. [44] developed convergence results for the multiobjective proximal gradient method. Additionally, Lapucci [31] studied the complexity of a wide class of multiobjective descent methods with nonconvex assumption. However, Chen et al. [8] noted that both theoretical and empirical results indicate that existing multiobjective first-order methods exhibit slow convergence due to objective imbalances. To address this challenge, Chen et al. [7] proposed a Barzilai-Borwein descent method for MOPs (BBDMO) that dynamically tunes gradient magnitudes using Barzilai-Borwein’s rule [3] in direction-finding subproblem. From a theoretical perspective, an improved linear convergence rate is confirmed by Chen et al. [8], demonstrating that Barzilai-Borwein descent methods can effectively mitigate objective imbalances. Despite the extensive study of the complexity analysis of descent methods for MOPs, corresponding results for VOPs have received little attention. As described by Chen et al. [6], the linear convergence rates of first-order descent methods for VOPs are influenced by $C$, the base of $K^{*}$ in direction-finding subproblem. This naturally leads to the following question:
@@
(Q2) Is $\Delta_{m}$ a good choice of base for +$\mathbb{R}^{m}_{+}$ in SDMO? If not, how to choose a better one?
The majorization-minimization principle is a versatile tool for designing novel algorithms, which has been successful employed in nonconvex optimization [30], constrained optimization [29], and incremental optimization [34, 27]. The core idea behind the majorization-minimization method is to iteratively minimize a sequence of easily solvable upper bounds of the objective function. In this paper, we extend majorization-minimization method to vector optimization, addressing the aforementioned questions. Our work aims to provide a unified framework and convergence analysis of descent methods for VOPs from a majorization-minimization perspective. The primary contributions of this paper are summarized as follows:
(i) Before presenting the majorization-minimization method and its convergence analysis, we first define the concepts of strong convexity and smoothness for a vector-valued function with respect to a partial order. Leveraging these properties, we extend the notion of the condition number to VOPs, which plays a pivotal role in establishing the linear convergence of first-order methods for VOPs.
@@
(ii) We devise a generic majorization-minimization descent method for VOPs and develop its convergence analysis. By selecting different surrogate functions, the general method can be reduced to several existing descent methods. We emphasize that the gap between the surrogate and objective functions significantly affects the performance of descent methods, which plays a central role in majorization-minimization optimization. Specifically, the steepest descent method for VOPs exhibits slow convergence due to the large gap between the surrogate and objective functions. To address this issue, we develop an improved descent method with a tighter surrogate function, resulting in improved linear convergence, and the rate of convergence is determined by the condition number. Interestingly, we show that selecting a tighter surrogate function is equivalent to using an appropriate base in the direction-find subproblem (see Remark 4.4). This provides a positive answer to (Q1). Especially, theoretical results suggest a tighter surrogate function by using Barzilai-Borwein method.
(iii) We propose a Barzilai-Borwein descent method for VOPs (BBDVO) with polyhedral cones. By reformulating the subproblem, we observe that BBDMO is essentially the SDMO with an appropriately chosen base in the direction-finding subproblem (see Remark 6.3). This observation provides a positive answer to (Q2) and offers new insights into BBDMO from a majorization-minimization perspective. Notably, a VOP with a polyhedral cone can be transformed into an MOP using a transform matrix, which is often used to define a specific polyhedral cone. We demonstrate that the performance of BBDVO is insensitive to the choice of the transform matrix, whereas the steepest descent method is highly sensitive to it. Numerical results also confirm that BBDVO is an efficient method for solving VOPs with polyhedral cones.
(iv) The fast convergence of Newton method for MOPs [16] comes at a computationally expensive per-step cost. By selecting an appropriate base and reformulating the subproblem in the Newton method, we propose an efficient second-order method for MOPs that coincides with the preconditioned Barzilai-Borwein method for MOPs (PBBMO) [9]. This formulation offers novel insights into PBBMO and highlights the indispensability of the Barzilai-Borwein rule in the quasi-Newton method from a majorization-minimization perspective.
@@
The remainder of this paper is organized as follows. Section 2 introduces the necessary notations and definitions for later use. In Section 3, we present a generic majorization-minimization descent method for VOPs and analyze its convergence rates under various convexity assumptions. Sections 4 and 5 explore the connections between different descent methods from the perspective of majorization-minimization. Section 6 proposes a Barzilai-Borwein descent method specifically designed for VOPs with a polyhedral cone. In Section 7, we study second-order methods from a majorization-minimization perspective. Section 8 provides numerical results to demonstrate the efficiency of the BBDVO. Finally, conclusions are presented at the end of the paper.
@@ 2.1 Vector optimization
By the continuous differentiability of $F(\cdot)$, $K$-convexity of $F(\cdot)$ is equivalent to
@@ 2.2 Strong convexity and smoothness
Comparing with the definition in [13], Definition 2.6 includes the case $\bm{\mu}\in{\rm{bd}}(K)$, then it reduces $K$-convexity when $\bm{\mu}=0$. Furthermore, tthe final statement in the definition establishes the uniqueness of the parameter $\bm{\mu}$, which is essential for the convergence analysis.
Lemma 2.2.
[25] Assume that the objective function $F(\cdot)$ is strongly $K$-convex with $\bm{\mu}\in{\rm int}(K)$, then $x^{*}\in\mathbb{R}^{n}$ is a $K$-stationary point of (VOP) if and only if $x^{*}$ is an efficient solution of (VOP).
By the continuous differentiability of $F(\cdot)$, strong $K$-convexity of $F(\cdot)$ is equivalent to
@@
In other words, the condition number of $F(\cdot)$ is determined only by $K$, not $C$.
In the following, we will show that the condition number can be reduced to that of MOPs [8].
Proposition 2.1.
For any $\bm{\ell},\bm{\mu}\in\mathbb{R}^{m}_{++}$, we have
$\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}=\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
Proof.
Without loss of generality, we assume that
$\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}\geq\frac{\bm{\ell}_{2}}{\bm{\mu}_{2}}\geq\cdots\geq\frac{\bm{\ell}_{m}}{\bm{\mu}_{m}}.$ (2)
Then, we infer that
$\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}\geq\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}=\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}.$
In the following, we need to prove
$\max\limits_{i\in[m]}\frac{\bm{\ell}_{i}}{\bm{\mu}_{i}}\geq\max\limits_{\lambda\in\Delta_{m}}\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}.$
It is sufficient to prove that
$\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}\geq\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}$
holds for all $\lambda\in\Delta_{m}.$ Let us first state the fact that for any $a_{1}/b_{1}\geq a_{2}/b_{2}>0,c,d>0$, we have
$\frac{a_{1}}{b_{1}}\geq\frac{ca_{1}+da_{2}}{cb_{1}+db_{2}}\geq\frac{a_{2}}{b_{2}}.$ (3)
For any $\lambda\in\Delta_{m}$, we denote $\mathcal{I}(\lambda):=\{i:\lambda_{i}>0\}$ with element $i_{1}<i_{2}<\cdots<i_{|\mathcal{I}(\lambda)|}$. Substituting $a_{1}/b_{1}={\bm{\ell}_{i_{1}}}/{\bm{\mu}_{i_{1}}}$, $a_{2}/b_{2}={\bm{\ell}_{i_{2}}}/{\bm{\mu}_{i_{2}}}$, $c=\lambda_{i_{1}}$ and $d=\lambda_{i_{2}}$ into (3) yields
$\frac{\bm{\ell}_{i_{1}}}{\bm{\mu}_{i_{1}}}\geq\frac{\lambda_{i_{1}}\bm{\ell}_{i_{1}}+\lambda_{i_{2}}\bm{\ell}_{i_{2}}}{\lambda_{i_{1}}\bm{\mu}_{i_{1}}+\lambda_{i_{2}}\bm{\mu}_{i_{2}}}\geq\frac{\bm{\ell}_{i_{2}}}{\bm{\mu}_{i_{2}}}.$
For $2\leq j\leq|\mathcal{I}(\lambda)|$, by substituting $a_{1}/b_{1}={\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\ell}_{i_{s}}}/{\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\mu}_{i_{s}}}$, $a_{2}/b_{2}={\bm{\ell}_{i_{j}}}/{\bm{\mu}_{i_{j}}}$, $c=1$ and $d=\lambda_{i_{j}}$ into (3) recursively, we have
$\frac{\bm{\ell}_{i_{1}}}{\bm{\mu}_{i_{1}}}\geq\cdots\geq\frac{\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\ell}_{i_{s}}}{\sum_{s\in[j-1]}\lambda_{i_{s}}\bm{\mu}_{i_{s}}}\geq\cdots\geq\frac{\sum_{s\in[|\mathcal{I}(\lambda)|]}\lambda_{i_{s}}\bm{\ell}_{i_{s}}}{\sum_{s\in[|\mathcal{I}(\lambda)|]}\lambda_{i_{s}}\bm{\mu}_{i_{s}}}.$
This, together with the definition of $\mathcal{I}(\lambda)$ and the relation (2), implies
$\frac{\bm{\ell}_{1}}{\bm{\mu}_{1}}\geq\frac{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}.$
Since $\lambda$ is chosen arbitrary, the proof is complete. ∎
Remark 2.6.
If $K=\mathbb{R}^{m}_{+}$, then $\kappa_{F,\preceq_{K}}=\max_{\lambda\in\Delta_{m}}{\sum_{i\in[m]}\lambda_{i}\bm{\ell}_{i}}/{\sum_{i\in[m]}\lambda_{i}\bm{\mu}_{i}}$. By Proposition 2.1, it follows that $\kappa_{F,\preceq_{K}}=\max_{i\in[m]}{\bm{\ell}_{i}}/{\bm{\mu}_{i}}$. The condition number reduces to that of MOPs [8].
@@
Remark 2.7.
Proposition 2.2 demonstrates that for a given vector-valued function, enlarging the order cone simplifies the corresponding vector optimization problem.
@@ 2       Choose a strongly $K$-convex surrogate function $G_{k}(\cdot)$ of $F(\cdot)-F(x^{k})$ near $x^{k}$
3       Update $x^{k+1}:=\mathop{\arg\min}_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x)}\right\rangle$
@@
    $F(x^{k+1})-F(x^{k})\preceq_{K}G_{k}(x^{k+1})$, where $x^{k+1}$ is the minimizer of $\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x)}\right\rangle$, furthermore, when $F(\cdot)-F(x^{k})\preceq_{K}G_{k}(\cdot)$ for all $x\in\mathbb{R}^{n}$, we call $G_{k}(\cdot)$ a majorizing surrogate;
@@
Let $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $x^{k+1}$ be the minimizer of $\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x)}\right\rangle$. Then, for all $x\in\mathbb{R}^{n}$, we have
@@
    $\left\langle{c^{*}_{k},F(x^{k+1})}\right\rangle+\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle\leq\left\langle{c^{*}_{k},F(x)}\right\rangle+\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle$, where $c^{*}_{k}$ is a maximizer of $\max_{c^{*}\in C_{k}}\min_{x\in\mathbb{R}^{n}}\left\langle{c^{*},G_{k}(x)}\right\rangle$.
@@
Assertion (i) directly follows by the $K$-smoothness of $H_{k}(\cdot)$ and the facts that $H_{k}(x^{k})=0$ and $JH_{k}(x^{k})=0$. Next, we prove the assertion (ii). By Sion’s minimax theorem [42], by denoting $c^{*}_{k}$ a maximizer of $\max_{c^{*}\in C_{k}}\min_{x\in\mathbb{R}^{n}}\left\langle{c^{*},G_{k}(x)}\right\rangle$, and $x^{k+1}$ the minimizer of $\min_{x\in\mathbb{R}^{n}}\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x)}\right\rangle$, we have $JG_{k}(x^{k+1})^{T}c^{*}_{k}=0$. This, together with the strong $K$-convexity of $G_{k}(\cdot)$, implies that